OpenStudy (rmrjr22):

1. (Molecular Kinetic Theory) Nitrogen gas is contained in a vessel; the mass of a nitrogen molecule is 4.68x10-26 kg. Every second 5.0x1023 nitrogen molecules strike a wall of area 8.0 cm2. If the molecules move at the rms speed of 300 m/s and strike the wall in a head-on perfectly elastic collision, find the pressure exerted on the wall.

OpenStudy (mos1635):

 P=F'/A P=(-F)/A P=N*(-ΔP/Δt)*1/A P=N*(-Δ(m*v))/Δt *1/A P=(N/Δt)*m*(-Δv)*(1/A) (N/Δt)=Every second 5.0x1023 nitrogen Δv=-2*v